Can you help me with this problem?
Find three consecutive integers whose sum is 321

Answer by checkley71(8403)

x+(x+1)+(x+2)=321
3x+3=321
3x=321-3
3x=318
x=318/3
x=106 for the first integer.
106+1=107 for the second
106+2=108 for the third
proof
106+107+108=321
321=321